OpenStudy (anonymous):

Solve –3(2x – 1) = x + 10. You must show all work.

OpenStudy (anonymous):

@Kristy143SIO

OpenStudy (anonymous):

Distribute Law: -6x + 3 = x + 10 -7x = 7 x=-1

OpenStudy (anonymous):

can u help me solve it but show all the work

OpenStudy (anonymous):

where u get -7 from

OpenStudy (anonymous):

@Yahoo!